# Mapping the Unseen: Unified Promptable Panoptic Mapping with Dynamic Labeling using Foundation Models

### v2

@@ 1 Introduction
In this paper, we present a prompt-based panoptic mapping pipeline that uses on-demand label generation based on natural language inputs. As a result of this approach, robots can dynamically acquire and apply object labels, allowing deeper semantic understanding and enhanced flexibility. We define Dynamic Labeling as the process of assigning semantically unified categories (see fig. 3) to detected objects in previously unseen environments, all while preserving the rich labels generated through open-vocabulary methods. Thorough evaluations and ablation studies on real-world and simulated datasets show that UPPM scene reconstruction is accurate and the dynamic label generation works well. This study opens the door to more natural human-robot communication and flexible machine perception in dynamic contexts. (Webpage)
@@ 3 Methodology
Figure 2: System Overview of Unified Promptable Panoptic Mapping (UPPM). The UPPM pipeline requires posed RGBD images. VLFE++ generates an open-vocabulary set of classes based on RGB inputs. For each class in the open-vocabulary set, STS predicts its size and parent class. This data is then fed into the Open-set Object Detection & Segmentation modules, where detection is refined by NMS to ensure enhanced accuracy. Combining outputs from previous steps with RGBD inputs, we adapt the panoptic mapping approach to accept dynamic labels (i.e. unified semantics), ensuring compatibility without overstating deviations. The final output is a dynamically labeled multi-resolution multi-TSDF map.
@@ 3.3 Open-set Object Detection and Segmentation
Object detection and segmentation are critical tasks in computer vision, particularly when dealing with unconstrained environments [21, 29, 30, 31]. Our method employs Grounding-DINO [20] as the object detector, querying it with curated textual labels for efficient semantic information extraction and object detection. Grounding-DINO generates bounding boxes, which are then used to prompt the Segment Anything Model (SAM) [13] to create high-quality instance masks for each detected object.
To further refine our approach, we employ Non-Maximum Suppression. During this stage, we address potential redundancies or errors introduced during object detection. Specifically, we focus on rectifying instances involving overlapping bounding boxes assigned identical labels or similar bounding box sizes but different labels, as shown in fig. 2. This approach distinct from per-class non-maximum suppression (NMS), which is commonly used in object detection [32]. In mathematical terms:
@@
[…]
@@
    In case of using the identified tags, when encounter bounding boxes that have equivalent areas but distinct class labels, the algorithm assigns the priority to the labels originating from captions rather than those derived from previously identified tags. As a result, only one bounding box will be used while others get removed through suppression.
    Mathematically, given two bounding boxes $b_{c}$ $b_{d}$ sharing same dimensions but disparate labels $y(b_{c})\neq y(b_{d})$, we prioritize retaining $b_{c}$ whenever its associated label stems from the caption. Otherwise, $b_{d}$ will be retained.
@@ 3.4 Unified Panoptic Mapping
The inputs of this component are the posed RGBD data, the open-vocabulary classes with their predicted sizes, the detected objects and panoptic segmentation.
In our study, we build upon the object-centric mapping framework introduced in [12]. So each submap contains all related data such as panoptic, instance, and class labels, as well as transformation and tracking data. Inspired by their semantically consistent representation of the environment, we adopt their notion of dividing the world into submaps for accurate temporal alignment and reduced computational complexity. However, we aim to improve the quality of objects contained within the maps. By integrating posed RGBD data with Unified Semantics, we ensure consistent object identification across different semantic labels through dynamic labeling, as illustrated in fig. 3. This approach simplifies classification and enhances the intuitive interaction with spatial information.
At the end of this stage, we will have a dynamically labeled multi-resolution multi-TSDF map.
@@ 4 Experiments
To evaluate our proposed method, we compare UPPM performance with MaskDino [5] and the reference data segmentation provided by the evaluation dataset. MaskDINO is a closed-set model that recognizes a wide range of categories, namely LVIS1 [33] categories (1203 categories). Our goal is to demonstrate the effectiveness of UPPM with the dynamically generated customized categories in comparison with these strong baselines. We conduct our experiments on three indoor evaluation datasets, one of them is simulated and the others are real-world datasets.
@@ 4.1 Experimental Setup
We measure the performance using root-mean-square error (RMSE), mean-absolute error (MAE), Chamfer distance, and coverage. As shown in eq. 2, the Chamfer distance $d_{\text{ch}}(G,R)$ is computed by combining the sum of RMSE in both directions, $\text{RMSE}_{G\to R}$ and $\text{RMSE}_{R\to G}$, providing the errors from the ground truth point cloud $G$ to the evaluated point cloud $R$ and vice versa.
@@
The experiments were conducted on a system with a 2.6 GHz CPU, an NVIDIA GeForce RTX 2060 6 GB GPU, and 16 GB of RAM, supplemented by a server featuring an NVIDIA GeForce RTX 2080 Ti 16 GB GPU and 32 GB of RAM.
@@ 4.2 Evaluations on the Flat dataset
In table 1, we present comparative quantitative results demonstrating the superior accuracy and competitive coverage of our UPPM model in comparison to both the dataset segmentation and the MaskDINO method [5].
@@
Method
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
Chamfer
dist.
[m]($\downarrow$)
Cover
-age
[%]($\uparrow$)
Dataset
Seg. [12]
1.27 2.24 0.66 0.80 3.05 71.30
MaskDINO
[5]
1.26 2.23 0.68 0.85 3.08 71.69
UPPM (Ours)
1.25 2.06 0.65 0.81 2.87 70.76
@@
It is worth noting that while MaskDINO achieves better coverage, a comprehensive comparison requires examining fig. 4. Upon inspection, it becomes apparent that MaskDINO has better reconstruction for structures like walls and floors. However, it fails to capture intricate details, like the paintings on the wall, which hold more semantic value. Conversely, UPPM successfully reconstructs these fine-grained elements.
@@ 4.3 Evaluations on ScanNet dataset
As demonstrated in table 2, UPPM exhibits better accuracy than MaskDINO [5] with a slight improvement in coverage. Although UPPM demonstrates strong performance, there is still room for improvement to surpass the accuracy scores achieved by the dataset segmentation [34].
@@ 4.4 Evaluations on the RIO dataset
The table 3 summarizes our quantitative findings, whereas fig. 5 shows our qualitative comparison. UPPM outperforms dataset segmentation [12] and MaskDINO [5] in accuracy while preserving good coverage.
@@ 5.2 Non-Maximum Suppression
To evaluate NMS (section 3.3) in our pipeline, we run the experiment on the Flat dataset. As shown in table 7, NMS improves coverage by $8.27\%$ gain, which proves that NMS widens the reconstructed area by eliminating overlapping or redundant predictions and enabling the pipeline to capture a wider range of discrete regions in the scene. This causes a small rise in Chamfer distance. The negligable decrease in point-to-point accuracy ($0.008[m]$ about $0.3\%$) can be attributed to the addition of less confident predictions in newly covered areas, and the occasional removal of valid overlapping predictions. In conclusion, NMS have significantly improved the overall performance.

### v3

@@ 1 Introduction
In this paper, we present a prompt-based panoptic mapping pipeline that uses on-demand label generation based on natural language inputs. As a result of this approach, robots can dynamically acquire and apply object labels, allowing deeper semantic understanding and enhanced flexibility. We define Dynamic Labeling as the process of assigning semantically unified categories (see fig. 3) to detected objects in previously unseen environments, all while preserving the rich labels generated through open-vocabulary methods. Thorough evaluations and ablation studies on real-world and simulated datasets show that UPPM scene reconstruction is accurate and the dynamic label generation works well at a rate of 0.65 FPS. This study opens the door to more natural human-robot communication and flexible machine perception in dynamic contexts. (Webpage)
@@ 3 Methodology
Figure 2: System Overview of Unified Promptable Panoptic Mapping (UPPM). The UPPM pipeline requires posed RGBD images. VLFE++ generates an open-vocabulary set of classes based on RGB inputs. For each class in the open-vocabulary set, STS predicts its size and parent class. This data is then fed into the Open-set Object Detection & Segmentation modules, where detection is refined by a custom NMS technique to ensure enhanced accuracy. Combining outputs from previous steps with RGBD inputs, we adapt the panoptic mapping approach to accept dynamic labels (i.e. unified semantics), ensuring compatibility without overstating deviations. The final output is a dynamically labeled multi-resolution multi-TSDF map.
@@ 3.3 Open-set Object Detection and Segmentation
Object detection and segmentation are critical tasks in computer vision, particularly when dealing with unconstrained environments [21, 29, 30, 31]. Our method employs Grounding-DINO [20] as the object detector, querying it with curated textual labels for efficient semantic information extraction and object detection. Grounding-DINO [20] generates bounding boxes, which are then used to prompt the Segment Anything Model (SAM) [13] to create high-quality instance masks for each detected object.
To further refine our approach, we address the potential redundancies and the limitations introduced during object detection [20] . For this we implement a custom Non-Maximum Suppression (NMS) technique to satisfy our needs that are not met by the traditional NMS [32], as shown in table 7. Traditional NMS, such as per-class NMS [32], typically removes redundant bounding boxes based on the Intersection-over-Union (IoU) and confidence scores for each detected object within the same class. However, our approach is described as follows:
@@
    Handling bounding boxes with similar sizes but different labels: In contrast to traditional per-class NMS [32], our approach resolves ambiguities when overlapping bounding boxes of the same size are assigned distinct class labels, as shown in fig. 2. For these cases, when the identified tags are used, we prioritize bounding boxes with labels generated from captions over those from the identified tags generated by the tagging head [18]. The decision to use captions is based on the assumption that captions provide more accurate contextual information over the tagging guidance [18]. Mathematically, given two bounding boxes $b_{c}$ and $b_{d}$ with similar dimensions but different labels $y(b_{c})\neq y(b_{d})$, we retain $b_{c}$ if its label comes from the caption; otherwise, $b_{d}$ is kept. A context-aware suppression approach is thus enabled, which sets it apart from typical NMS systems that may rely purely on confidence ratings and IoU levels.
@@
    Handling overlapping boxes with identical labels: Let Bi denote the set of all bounding boxes extracted from the object detector, such that $|B_{i}|\geqslant 1$. For any pair ($b_{a},b_{b}$), where $b_{a},b_{b}\in B_{i}$, if their Intersection-over-Union (IoU) exceeds a predefined threshold IoU($b_{a},b_{b}$) $\geqslant\tau$, and share the exact class label y($b_{a}$) = y($b_{b}$), one of the duplicates will be suppressed based on criteria like confidence scores, as shown in fig. 2.
@@ 3.4 Unified Panoptic Mapping
The inputs of this component are the posed RGBD data, the open-vocabulary classes with their predicted sizes, the detected objects and panoptic segmentation. This component aims to create a dynamically-labeled multi-resolution volumetric map that maintains semantic consistency and efficiently manages spatial and temporal data.
In our study, we extend the the object-centric mapping framework introduced in [12], which structures the environment into submaps. Each submap contains geometric and semantic data, such as panoptic, instance, and class labels, as well as transformation and tracking data. This submap-based structure enables efficient processing of large-scale environments by dividing the world into smaller, manageable parts, significantly reducing computational complexity. The novelty of our work lies in combining the strengths of both open-vocabulary and closed-vocabulary approaches by enriching submap objects with open-vocabulary classes, while simultaneously assigning semantically unified categories to detected objects in previously unseen environments. By integrating posed RGB-D data with unified semantics, we ensure consistent object identification across different semantic labels through dynamic labeling, as illustrated in fig. 3. This approach simplifies classification and enhances intuitive interaction with spatial information.
At the end of this stage, the Unified Panoptic Mapping component produces a dynamically labeled multi-resolution multi-TSDF map. Each object in this map is not only accurately localized but also semantically labeled, allowing for high-quality 3D reconstruction and facilitating downstream tasks such as navigation and localization.
@@ 4 Experiments
To evaluate our proposed method, we compare UPPM performance with MaskDINO[5] , OpenSeeD [33], OMG-Seg [34] and the reference data segmentation provided by the evaluation dataset. MaskDINO is a closed-set model that recognizes a wide range of categories, namely LVIS1 [35] categories (1203 categories). While OpenSeeD [33], and OMG-Seg [34] are open-vocabulary segmentation models. Our goal is to demonstrate the effectiveness of UPPM with the dynamically generated custom categories in comparison with these strong baselines. We conduct our experiments on three indoor evaluation datasets, one of them is simulated and the others are real-world datasets.
@@ 4.1 Experimental Setup
We measure the performance using root-mean-square error (RMSE), mean-absolute error (MAE), Chamfer distance, F1-Score, and coverage. As shown in eq. 2, the Chamfer distance $d_{\text{ch}}(G,R)$ is computed by combining the sum of RMSE in both directions, $\text{RMSE}_{G\to R}$ and $\text{RMSE}_{R\to G}$, providing the errors from the ground truth point cloud $G$ to the evaluated point cloud $R$ and vice versa.
@@
To measure the ability to reconstruct the fine-grained details in the scene, we utilize F1-Score after removing the background classes like floor, ceiling, and wall. For each point, we consider the reconstructed value correct if its absolute difference from the ground truth is within a specified threshold. The F1-score is then computed, providing a single, balanced measure of reconstruction accuracy that considers both false positives and false negatives.
The experiments were conducted on a system with a 2.6 GHz CPU, an NVIDIA GeForce RTX 2060 6 GB GPU, and 16 GB of RAM, supplemented by a server featuring an NVIDIA GeForce RTX 2080 Ti 11 GB GPU and 32 GB of RAM.
@@ 4.2 Evaluations on the Flat dataset
In table 1, we present comparative quantitative results demonstrating the superior accuracy and competitive coverage of our UPPM model in comparison to both the dataset segmentation and the closed-vocabulary MaskDINO method [5]. On the other hand, UPPM shows the best coverage, and competent chamfer distance in comparison with the open-vocabulary approaches [33, 34].
@@
Method MAE RMSE MAE RMSE Chamf- F1- Cov. FPS
[…]
Dataset Seg. [12] 1.27 2.24 0.66 0.80 3.05 - 71.30 -
MaskDINO [5] 1.26 2.23 0.68 0.85 3.08 76.01 71.69 1.7
OpenSeeD [33] 1.18 1.91 0.653 0.8 2.71 76.43 66.5 0.41
OMG-Seg [34] 1.14 1.81 0.67 0.81 2.62 77.99 68.03 0.38
UPPM (Ours) 1.25 2.06 0.65 0.81 2.87 79.54 70.76 0.65
@@
It is worth noting that while MaskDINO achieves better coverage, a comprehensive comparison requires examining fig. 4. Upon inspection, it becomes apparent that MaskDINO has better reconstruction for structures like walls and floors. However, it fails to capture intricate details, like the paintings on the wall, which hold more semantic value. Conversely, UPPM successfully reconstructs these fine-grained elements. This qualitative observation is backed by the F1-score results. As shown in table 1, we calculate the metric for both the closed- and open-vocabulary methods after removing the background classes such as wall, floor, and ceiling to have more focus on the intricate details of the scene. The F1-score results show the ability of UPPM to reconstruct the fine-grained details compared with the other methods.
As shown in table 1, we report the frames-per-second (fps) tested on an RTX 2080 Ti NVIDIA GPU for our pipeline by taking the average computing time with batch size 1 on the flat dataset.
@@ 4.3 Evaluations on ScanNet dataset
As demonstrated in table 2, UPPM exhibits better accuracy and slightly better coverage in comparison with the other state-of-the-art methods [5, 33, 34]. Although UPPM demonstrates strong performance, there is still room for improvement to surpass the accuracy scores achieved by the dataset segmentation [36].
@@ 4.4 Evaluations on the RIO dataset
The table 3 summarizes our quantitative findings, whereas fig. 5 shows our qualitative comparison. UPPM have the best accuracy while preserving a good coverage.
@@ 5.2 Non-Maximum Suppression
To show the strengths of the custom NMS over the traditional NMS in our pipeline, We run the pipeline once with the traditional NMS and then with our custom NMS. Upon manual inspection, our custom NMS appeared to detect possible duplicates better compared to the traditional NMS as the results shown in table 7.
Table 7: NMS [32] vs. our custom NMS on the Flat dataset [12].
Method Precision [%]($\uparrow$) Recall [%]($\uparrow$) F1-Score [%]($\uparrow$)
NMS [32] 91.3 63.64 75.0
Custom NMS (Ours) 94.6 100.0 97.22
To evaluate how the custom NMS (section 3.3) affect our pipeline, we run the experiment on the Flat dataset. As shown in table 8, NMS improves coverage by $8.27\%$ gain, which proves that NMS widens the reconstructed area by eliminating overlapping or redundant predictions and enabling the pipeline to capture a wider range of discrete regions in the scene. This causes a small rise in Chamfer distance. The negligible decrease in point-to-point accuracy ($0.008[m]$ about $0.3\%$) can be attributed to the addition of less confident predictions in newly covered areas, and the occasional removal of valid overlapping predictions. In conclusion, NMS have significantly improved the overall performance.